A car moving on straight road accelerates from speed of 4.1m/s to a speed of 6.9m/s in 5.0s
Then average acceleration is ?

Grade:11

2 Answers

Arun
5 years ago
 
Initial velocity,u=4.1m/s
Final velocity ,v=6.9m/s
Time taken=5.0secs
We know, acceleration=(v-u)/t where t=time.
Therefore, acceleration=(6.9-4.1)/5 m/sec square
                                     =2.8/5 m/sec square
                                     =0.56 m/sec square(ANS)
 
 
Regards
arun
kotte devika
2 years ago
given,
Initial velocity(u)=4.1m/s
Final velocity (v)=6.9m/s
Time(t)=5.0s
W.K.T..,
acceleration=(v-u)/t  a=(6.9-4.1)/5 m/s2
                                     =2.8/5 m/s2
                                     =0.56 m/s2
  Therefore a=0.56m/s2